# Field-based Molecule Generation

### v1

@@ Abstract
This work introduces FMG, a field-based model for drug-like molecule generation. We show how the flexibility of this method provides crucial advantages over the prevalent, point-cloud based methods, and achieves competitive molecular stability generation. We tackle optical isomerism (enantiomers), a previously omitted molecular property that is crucial for drug safety and effectiveness, and thus account for all molecular geometry aspects. We demonstrate how previous methods are invariant to a group of transformations that includes enantiomer pairs, leading them invariant to the molecular R and S configurations, while our field-based generative model captures this property.
@@ 1 Introduction
Developing generative models for a specific data modality requires carefully choosing the generative method, feature representation, and neural network (NN) architecture. Field-based representations offer a unifying perspective since modeling continuous domain functions is needed in many domains. In this work, we show how field-based molecule representations can be modeled successfully, achieving competitive performance, compared to the point-cloud molecule representation counterpart.
Organic molecules generally contain a restricted number of atom types, bound mostly through covalent bonds. Most atoms in organic molecules are non-ionized, and distances between two bonded atoms should follow empirically attested values. In an attempt to achieve realistic bond distances, rotation equivariant NN architectures have been devised (Peng et al., 2023b; Guan et al., 2023). The developed methods only receive pairwise atom distances, which are rotation invariant.
@@
While the aforementioned properties clearly incentivize the development of rotational equivariant architectures, this may come at the cost of potentially incorrect enantiomer generation. Enantiomers are mirror-image pairs of atoms, where one of the molecules (chiral configurations) cannot be superimposed to the other through any rotation and translation (see Fig 1). They are present in any molecule with a tetrahedral carbon configuration with four unique groups. In the popular drug-like dataset GEOM, about 27% of the molecules contain such chiral configurations, and they can be crucial in the potency and safety of drugs. In ibuprofen, albuterol, and naproxen, one of the two enantiomers was shown to be more effective, while in thalidomide and propoxyphene, the incorrect enantiomer is toxic and can lead to severe side effects (Smith, 2009).
Most prior work that generates molecular conformations focuses on methods that are invariant to the special orthogonal group of transformations in three dimensions, denoted SO(3). This means that any rotation $R$ multiplied with a vector of atom coordinates $\mathbf{m}$ will leave their corresponding NN activations unchanged. In this work, we discuss and show why these methods are, in fact, invariant to any orthogonal group transformation, of which SO(3) is a subset, and how this leads to unwanted invariance wrt different chiral configurations of molecules.
Mirror-imaging a molecule may lead to different chiral molecules for certain tetrahedral configurations and can be achieved by multiplying its atom positions with a reflection matrix $R$, which is orthogonal and satisfies $\operatorname{det}R=-1$. Expressing pairwise differences as vectors $\mathbf{d}=\mathbf{x}_{i}-\mathbf{x}_{j}$, we note that Euclidean distance will therefore be invariant to reflections, since $||\mathbf{d}||_{2}=||R\mathbf{d}||_{2}$, for any orthogonal matrix $R$. Neural networks that only receive pairwise atom Euclidean distances will be unable to distinguish chiral configurations. We conduct experiments that also empirically attest to this fact, and illustrate how the problem cannot be trivially solved in the context of point-cloud formulations, even when removing the SO(3) invariance of the NN. Conversely, our field-based method perfectly captures the chiral distribution in the simplified chiral experiment we design.
We devise a three-dimensional multi field-based molecule representation, where atoms and bonds are represented as isotropic radial basis functions. Previous work using similar input representations (Pinheiro et al., 2023; Ragoza et al., 2020) has shown that working in this input space is challenging. Namely, the methods rely on evaluating the RBF fields at specific grid locations, creating very large tensors, compared to point-cloud representations. As such, GNN or transformer-based architectures are computationally prohibitive, and CNNs are the only viable option. While CNNs only have a translation invariance inductive bias, we show that defining simple, yet effective, reference rotations for molecules enables powerful U-Net architecture to achieve very good results, while preserving correct chirality.
@@
Figure 1: Relative bond angles and distances are not sufficient to represent chirality. Top: A chiral molecule can not be rotated to its mirror image. Bottom: Chiral S-ibuprofen is an efficient COX-inhibitor resulting in pain and inflammation relief, while the mirror R-ibuprofen is not (Evans, 2001).
@@ 2 Related works
A large number of generative models to represent molecules have been proposed, which we distinguish primarily by their molecular representation instead of generative method families.
@@
SMILES is a string-based representation in which molecular graphs can be sufficiently encoded (Weininger, 1988), and has been employed in autoregressive models (Segler et al., 2018) and variational autoencoder (VAE) based generative models (Gómez-Bombarelli et al., 2018; Kusner et al., 2017). There are a few limitations when using smiles representations directly. First, similar molecular graphs may determine very different smiles encodings (e.g., measured by Levenshtein distance). Second, there can be multiple different smile representations for the same molecule, and the atom ordering in the string does not reflect the true, physical distance between the atoms.
@@
An alternative molecule representation consists of molecular graphs, where vertices encode atom type information, and their covalent bonds are represented through adjacency matrices. VAEs (Liu et al., 2018; Jin et al., 2018), flow-based models (Shi et al., 2020), and reinforcement learning (Shi et al., 2020; You et al., 2018; Zhou et al., 2019) methods were developed. The main drawback of only generating molecular graphs is that the stereochemical properties of molecules cannot be captured, and the functionality of various drug-like molecules depends on their three-dimensional conformation.
@@
The three-dimensional configuration of a molecule can be encoded as $\mathbb{R}^{3}$ point clouds of atoms. Molecules reside in the $\mathbb{R}^{3}$ space, and their conformations are symmetric w.r.t. global translations and rotations. Inspired by this, methods that adhere to the symmetries of the Euclidean group in three dimensions (E(3) equivariant) have been developed. Autoregressive models (Gebauer et al., 2018, 2019) have been developed, with the notable downside that they introduce an artificial atom order. GNN parametrized denoising diffusion models (Guan et al., 2023; Peng et al., 2023b; Vignac et al., 2023; Xu et al., 2023; Wu et al., 2022) have also been employed, which, compared to their autoregressive counterparts, remove the arbitrary order of atom generation. Although rotational invariance is unquestionably well motivated, these methods also encode the undesired reflection invariance, which is part of the E(3) symmetries.
@@
The $\mathbb{R}^{3}$ coordinates of atoms in molecules can be used as the mean positions of Gaussian or other densities, which fill the three-dimensional space. These densities can be interpreted as fields, where each molecule is represented by a function on $\mathbb{R}^{3}$, with values given by the densities of the atoms. Parameterizing the generative models with CNNs, we crucially distinguish this approach from the previously developed methods that use point cloud formulations by the E(3) symmetries they adhere to. Namely, our approach only respects translation invariance, while the point cloud approaches respect all E(3) symmetries. At the expense of losing the rotational invariance, this allows us to generate chiral conformations according to data statistics, which cannot be captured by reflection invariant models. This input representation has been previously modeled using VAE (Ragoza et al., 2020) and score-based models (Pinheiro et al., 2023), with some degree of success.
@@ 3.1 Field representation
We consider a molecule as a set of atoms $\{a_{i},\mathbf{m}_{i}\}$ and a set of bonds $\{b_{ij},\mathbf{m}_{ij}\}$, with atom types $a_{i}\in\mathcal{A}=\{\texttt{H},\texttt{C},\texttt{N},\texttt{O},\texttt{P},\texttt{S},\ldots\}$, bond types $b_{ij}\in\mathcal{B}=\{b_{1},b_{2},b_{3}\}$, and we assume atom locations $\mathbf{m}_{i}\in\mathbb{R}^{3}$ and mid-point bond locations $\mathbf{m}_{ij}=\frac{\mathbf{m}_{i}+\mathbf{m}_{j}}{2}$ in Ångstroms. We represent a molecule using a collection of continuous 3D fields, one for each atom and bond type. We define the field for atom type $a$, $u_{a}$, as a mixture of RBF functions positioned on the atoms’ centers
@@
where $N$ is the number of atoms, $\mathds{1}[\cdot]$ is the indicator functions, the variance $\sigma_{a}^{2}$ represents the size of atom type $a$, and $\gamma_{a}$ normalises the channels to unit interval $u_{a}(\mathbf{x})\in[0,1]$ (see Appendix B for field parameter values). We create bond fields similarly
@@
We evaluate a molecule’s atom and bond fields $u_{a},u_{b}$ at a fixed set $\mathcal{G}$ of $H\times W\times D$ locations, and create a vector-valued field based on which multi-channel 4D tensors are derived
@@
where $K=|\mathcal{A}|+|\mathcal{B}|$ is the number of channels. We use tensor $\mathbf{u}$ as the input molecule representation for the diffusion model (See Figure 2).
Figure 2: The hydrogen channel voxel grid values $\mathbf{u}_{\text{H}}$ s.t., $\mathbf{u}_{\text{H}}>0.1$, are visualized. All atom positions and bonds are depicted through circles and line plots, respectively, for reference.
Figure 3: The faster atom schedule in the forward process results in the model first generating the molecule structure through its bonds, followed by atoms.
@@ 3.2 Diffusion model
We consider a diffusion generative model to generate molecular fields $\mathbf{u}$. In diffusion models, the data is incrementally noised by a forward process, and our goal is to learn a denoising reverse process parameterised by $\theta$. We consider the field tensors $\mathbf{u}_{t}$ over time $t\in[0,1]$, where we assume the observed or generated tensors are $\mathbf{u}_{0}$ at time $t=0$.
We follow the denoising diffusion probabilistic model (DDPM) proposed in (Sohl-Dickstein et al., 2015; Ho et al., 2020b; Kingma et al., 2021), and refer the reader to the original publications for full derivations. We note that our tensor fields are effectively multi-channel 3D images, and thus image-based diffusion models translate to molecular voxel grid fields in a straightforward manner.
@@
In practice, it is often more efficient to train a neural network to predict the added noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ of
$\mathbf{u}_{t}=\alpha_{t}\mathbf{u}_{0}+\sigma_{t}\boldsymbol{\epsilon},$ (8)
than to predict the signal $\mathbf{u}_{0}$. We then train a neural network $\boldsymbol{\epsilon}_{\theta}(\mathbf{u}_{t})$ to predict the noise $\boldsymbol{\epsilon}$, and follow the above equations with a substitution
$\hat{\mathbf{u}}_{\theta}=\mathbf{u}_{t}/\alpha_{t}-\sigma_{t}\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{u}_{t},t)/\alpha_{t}.$ (9)
@@
The diffusion model admits a variational lower bound (Kingma et al., 2021), but earlier works have argued for a simplified loss (Ho et al., 2020b)
$\mathcal{L}_{\mathrm{simple}}$ $=\mathbb{E}_{\mathbf{u},t,\boldsymbol{\epsilon}}\Big{[}||\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{u}_{t},t)||^{2}\Big{]},$ (10)
where we learn to denoise arbitrary combinations of observations $\mathbf{u}$, noises $\boldsymbol{\epsilon}$ and noise scales $t\sim\mathrm{U}(0,1)$. We also opt for the simple loss $\mathcal{L}_{\mathrm{simple}}$.

### v2

# E(3)-equivariant models cannot learn chirality: Field-based molecular generation

@@ Abstract
Obtaining the desired effect of drugs is highly dependent on their molecular geometries. Thus, the current prevailing paradigm focuses on 3D point-cloud atom representations, utilizing graph neural network (GNN) parametrizations, with rotational symmetries baked in via E(3) invariant layers. We prove that such models must necessarily disregard chirality, a geometric property of the molecules that cannot be superimposed on their mirror image by rotation and translation. Chirality plays a key role in determining drug safety and potency. To address this glaring issue, we introduce a novel field-based representation, proposing reference rotations that replace rotational symmetry constraints. The proposed model captures all molecular geometries including chirality, while still achieving highly competitive performance with E(3)-based methods across standard benchmarking metrics. Code is available at https://dumitrescu-alexandru.github.io/FMG-web/.
@@ 1 Introduction
Developing generative models for specific data modalities requires carefully choosing the generative method, feature representation, and NN architecture. Field-based representations may offer a unifying perspective, extending the usage of highly successful architectures from natural language or images to other input modalities Zhuang et al. (2023). We develop Field-based Molecule Generation (FMG), a generative model of three-dimensional field representations of molecules, which encode the joint distribution of molecular graphs and 3D conformations.
Organic molecules generally contain a restricted number of atom types, bound mostly through covalent bonds, and they can be used as drugs, inhibiting or promoting biological processes. Their effectiveness crucially depends on successfully binding and interacting with proteins or other small molecules in the organism. In turn, the interaction between two molecules can only be determined once their 3D conformations are known. This motivated the development of many generative models for molecular conformations. Recently, the vast majority of work has been done on generative model parameterizations that are invariant to the Euclidean group of isometries in 3D, E(3) (Hoogeboom et al., 2022; Peng et al., 2023; Vignac et al., 2023b; Xu et al., 2023; Wu et al., 2022).
[…]
@@
Chiral centers are often tetrahedral carbon configurations containing four unique groups. In molecular biology, chirality is crucial to the potency and safety of drugs. In ibuprofen, albuterol, and naproxen, one of the two enantiomers was shown to be more effective, while in thalidomide and propoxyphene, the incorrect enantiomer is toxic and can lead to severe side effects (Smith, 2009). In proteins, out of 21 standard amino acids, only Glycine is not chiral. Although classically it was believed that only one chiral configuration is present in nature (Devínsky, 2021), patients with disorders such as cataracts, macular degeneration, arteriosclerosis, and Alzheimer’s disease present both (Masters et al., 1977; Fujii et al., 2018). In the drug-like molecule dataset, GEOM, 27% of the molecules contain chiral centers.
In essence, reflections are distance preserving, and therefore parameterizations that utilize Euclidean distances will be unable to distinguish enantiomers. Later on, we prove theoretical results that show how designing enantiomer (reflection) variant features (SE(3) invariant) grows the complexity from $\mathcal{O}(n^{2})$ (for pairwise Euclidean distances between $N$ atoms) to $\mathcal{O}(n^{4})$. To account for chiral centers, we instead chose to replace rotation symmetry constraints, and propose reference rotations, which bring our performance to highly competitive results with E(3) invariant methods.
We organize our contributions as follows:
Theoretical:
We prove that functions of E(3) invariant features fail to properly model chiral molecules. We then link the parametrizing function’s E(3) invariance to the resulting generative E(3) equivariance. Finally, we prove that SE(3)-invariant, chiral aware features can only be functions of a least four $\mathbb{R}^{3}$ vectors, and the number of such features scales in $\mathcal{O}(n^{4})$.
Practical:
We provide solutions to the challenges of handling 3D fields. We present a method of reorienting molecules to deterministic references, introduce bond fields, describe our diffusion process design, and complete our pipeline with molecular graph optimization based on fields.
Empirical:
Put together, our technical contributions result in SOTA or highly competitive performance compared to point-cloud approaches, while keeping chiral-awareness.
@@
Figure 1: E(3) invariant spatial features (e.g., relative bond angles and distances) are not sufficient to represent chirality. Subfigure a) depicts a general chiral molecule pair, which cannot be superimposed by rotations. b) Chiral S-ibuprofen is an efficient COX-inhibitor resulting in pain and inflammation relief, while the mirror R-ibuprofen is not (Evans, 2001).
@@ 2 Related works
A large number of generative models for molecules have been proposed. We distinguish them primarily by their molecular representation instead of generative method families.
@@
SMILES is a string-based representation in which molecular graphs can be sufficiently encoded (Weininger, 1988), and has been employed in autoregressive (Segler et al., 2018) and variational autoencoder (VAE) models (Gómez-Bombarelli et al., 2018; Kusner et al., 2017). An important limitation of SMILES used as input representations for machine learning is that similar molecular graphs may have very different SMILES encodings (Jin et al., 2018).
@@
Alternative molecule representations consist of molecular graphs, where vertices encode atom types and their covalent bonds are represented through adjacency matrices. VAEs (Liu et al., 2018; Jin et al., 2018), flow-based models (Shi et al., 2020; Verma et al., 2022), reinforcement learning (Shi et al., 2020; You et al., 2018; Zhou et al., 2019) and diffusion-based methods (Vignac et al., 2023a; Mercatali et al., 2024) have been developed. The main drawback of only generating molecular graphs is that full molecular geometry specification is missing, and the functionality of various drug-like molecules depends on it.
@@
The three-dimensional configuration of a molecule can be encoded as $\mathbb{R}^{3}$ point clouds of atoms. Molecules reside in the $\mathbb{R}^{3}$ space, and their conformations are symmetric w.r.t. global translations and rotations. Inspired by this, E(3) equivariant generative methods have been developed. Autoregressive models (Gebauer et al., 2018; 2019), GNN parametrized diffusion (Hoogeboom et al., 2022; Peng et al., 2023; Vignac et al., 2023b; Xu et al., 2023; Wu et al., 2022), and transformer parametrized diffusion (Hua et al., 2023; Huang et al., 2024) methods have recently gained a lot of interest. Although rotational invariance is well motivated, these methods also encode the undesired invariance to reflections, which are part of the E(3) group.
@@
The $\mathbb{R}^{3}$ coordinates of atoms in molecules can be used as the mean positions of e.g. Gaussians, which fill the three-dimensional space. These densities can be interpreted as fields, where each molecule is represented by a function on $\mathbb{R}^{3}$, with values given by the densities of the atoms. We crucially distinguish this approach from the previously developed methods that use point cloud formulations by the E(3) symmetries they adhere to. Namely, our approach only respects translation invariance through the CNN inductive biases, while point cloud approaches respect all E(3) symmetries. At the cost of losing rotational invariance, our method can generate chiral conformations according to data statistics, while E(3) invariant formulations cannot. Similar input representations have been previously modeled using VAE (Ragoza et al., 2020) and score-based models (Pinheiro et al., 2023). The latter is the closest to our work, which bases its method on the neural empirical Bayes framework and only generates atom voxels, based on which they use post-generation bond extraction. Using diffusion models, we show that field representations can reach SOTA performance while generating the full molecule specification (including bond fields).
Chiral-aware methods
Prior, chiral-aware discriminative methods for molecular analysis have been developed in (Adams et al., 2022; Liu et al., 2022). Jing et al. (2022) also accounts for chirality in a generative setting using a diffusion model, but considers the conditional generation of 3D molecular conformations given a molecular graph $p(C|G)$, while the scope of this work is on the joint generation $p(C,G)$. Finally, Luo & Ji (2022) uses (Liu et al., 2022) NN parametrization of a joint $p(C,G)$ AR generative model that, in principle, is chiral-aware, but does not discuss chirality. AR methods need to define arbitrary node order generation, and whether they can reach diffusion model performance despite this fact remains an open question. In general, chiral-aware SE(3) diffusion models on point clouds are surprisingly difficult to develop.
3 Chirality and isometry groups
Most recent work on molecule generation model E(3) equivariant distributions. Although molecules respect rotation and translation symmetries, chiral molecules are not symmetric to reflections. In such cases, the reflected version of a therapeutic drug can become toxic (Smith, 2009).
Proposition 1.
If $p_{\phi}$ is an E(3) invariant probability distribution, then $p_{\phi}(\mathbf{m})=p_{\phi}(\mathbf{m}^{\prime})$, where $\mathbf{m}$ and $\mathbf{m}^{\prime}$ are the $\mathbb{R}^{3\times N}$ positions of enatiomer pairs of molecules with $N$ atoms.
Proposition 1 is a consequence of the fact that $p_{\phi}(\mathbf{m})=p_{\phi}({\bm{R}}\mathbf{m}),\forall$ ${\bm{R}}$ orthogonal, $\det{\bm{R}}=\pm 1$. This implies that NN parameterizations of point-cloud diffusion models should be SE(3) invariant, but we argue that this becomes intractable.
Lemma 2.
Let $f:\mathbb{R}^{m\times n}\rightarrow\mathcal{S}$, for some set $\mathcal{S}$, be a function of $m$ $n$-dimensional vectors. If $m\leq n$ and $f$ is SE(n) invariant, then $f$ is $E(n)$ invariant.
For example, in $\mathbb{R}^{3}$, the reflection of a planar object ($m=n=3$) can be recovered by rotations and translations. This implies that there is no SE(3) invariant function that, applied on triplets of atom positions, would distinguish enantiomer pairs.
Proposition 3.
Let $n$ be the number of atoms of a molecule. The number of features that encode all chiral configurations of random point clouds scales in $O(n^{4})$.
The result implies that SE(3) (and not E(3)) invariant features scale prohibitively with the number of atoms.
Proofs can be found in Appendix B. There, we also concretely describe the model types we refer to in Proposition 3, which include all SOTA diffusion models (Hoogeboom et al., 2022; Peng et al., 2023; Vignac et al., 2023b; Xu et al., 2023; Wu et al., 2022).
@@ 4.1 Field representation
We consider a molecule as a set of atoms $\{a_{i},\mathbf{m}_{i}\}$ and a set of bonds $\{b_{ij},\mathbf{m}_{ij}\}$, with atom types $a_{i}\in\mathcal{A}=\{\texttt{H},\texttt{C},\texttt{N},\texttt{O},\texttt{P},\texttt{S},\ldots\}$, bond types $b_{ij}\in\mathcal{B}=\{b_{1},b_{2},b_{3}\}$, and we assume atom locations $\mathbf{m}_{i}\in\mathbb{R}^{3}$ and mid-point bond locations $\mathbf{m}_{ij}=\frac{\mathbf{m}_{i}+\mathbf{m}_{j}}{2}$ in Ångströms.
We start by representing a molecule using a collection of continuous 3D fields, one for each atom and bond type. We define the field for atom type $a$, $u_{a}$, as a mixture of radial basis functions (RBFs) positioned on the atoms’ centers
@@
where $N$ is the number of atoms, $\mathds{1}[\cdot]$ is the indicator function, the variance $\sigma_{a}^{2}$ represents the size of atom type $a$, and $\gamma_{a}$ normalises the channels to unit interval $u_{a}(\mathbf{x})\in[0,1]$ (see Appendix C for field parameter values). We create bond fields similarly
@@
We evaluate a molecule’s atom and bond fields $u_{a},u_{b}$ at a fixed set $\mathcal{G}$ of $H\times W\times D$ (height, width, and depth) locations, and create a vector-valued field based on which multi-channel 4D tensors are derived
@@
where $K=|\mathcal{A}|+|\mathcal{B}|$ is the number of channels. We use tensor $\mathbf{u}$ as the input molecule representation for the diffusion model.
Figure 2: a) Atom $\mathbf{u}_{a}$ and bond $\mathbf{u}_{b}$ fields noise schedules. b) Sampling illustration. A subset of 200 3D locations with the highest values are shown for all $\mathbf{u}_{a}$ and $\mathbf{u}_{b}$ channels. c) Extracting atoms and bonds from the resulting $\mathbf{u}_{a}$ and $\mathbf{u}_{b}$ fields at $t=0$. d) Visualizing the optimized atoms and bonds.
@@ 4.2 Diffusion model
We consider a diffusion generative model to generate molecular fields $\mathbf{u}$. In diffusion models, the data is incrementally noised by a forward process, and our goal is to learn a denoising reverse process parameterised by $\theta$. We consider the field tensors $\mathbf{u}_{t}$ over time $t\in[0,1]$, where $\mathbf{u}_{0}$ at time $t=0$ are the observed or generated tensors. Figure 2b depicts an example of our denoising reverse process.
We follow the denoising diffusion probabilistic model (DDPM) proposed in (Sohl-Dickstein et al., 2015; Ho et al., 2020; Kingma et al., 2021), and refer the reader to the original publications for full derivations. We note that our tensor fields are effectively multi-channel 3D images, and thus image-based diffusion models translate to molecular voxel grid fields in a straightforward manner.
@@
In practice, it is often more efficient to train a neural network to predict the added noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ of
$\mathbf{u}_{t}=\alpha_{t}\mathbf{u}_{0}+\sigma_{t}\bm{\epsilon},$ (8)
than to predict the signal $\mathbf{u}_{0}$. We then train a neural network $\bm{\epsilon}_{\theta}(\mathbf{u}_{t})$ to predict the noise $\bm{\epsilon}$, and follow the above equations with a substitution
$\hat{\mathbf{u}}_{\theta}=\mathbf{u}_{t}/\alpha_{t}-\sigma_{t}\hat{\bm{\epsilon}}_{\theta}(\mathbf{u}_{t},t)/\alpha_{t}.$ (9)
@@
The diffusion model admits a variational lower bound (Kingma et al., 2021), but earlier works have argued for a simplified loss (Ho et al., 2020)
$\mathcal{L}_{\mathrm{simple}}$ $=\mathbb{E}_{\mathbf{u}_{0},t,\bm{\epsilon}}\Big{[}||\bm{\epsilon}-\hat{\bm{\epsilon}}_{\theta}(\mathbf{u}_{t},t)||^{2}\Big{]},$ (10)
where we learn to denoise arbitrary combinations of observations $\mathbf{u}_{0}$, noises $\bm{\epsilon}$ and noise scales $t\sim\mathrm{U}(0,1)$. We also opt for the simple loss $\mathcal{L}_{\mathrm{simple}}$.